OpenStudy (anonymous):

Given the equation and solution, find the missing coefficient. 5x − 2 = -7x + Solution: x = 5

OpenStudy (anonymous):

5x-2=-7x+58 5x+7x=58+2 12x=60 x=5 so,the missing value is '58'...

OpenStudy (anonymous):

so,waltersa....is it right????